# Integrating Graceful Degradation and Recovery through Requirement-driven Adaptation

### 18 Jan 2024 v1

@@ IV Runtime Adaptation Architecture
There are three major components in the proposed adaptation architecture. First, the event detector looks for degradation or restoration events in the environment. Second, the requirement evaluator determines the achievable requirement based on the current environmental conditions. Third, the degradation and recovery planner plan future system actions based on the changing requirements. For example, if visibility decreases below a certain threshold and the weakened requirement to regain contact within the next 5 seconds, the degradation planner may institute a wait action instead of diving directly or diving with a lower speed.
When activated by the event detector, the requirement evaluator takes various inputs: (1) an environmental event, (2) the current system requirement, (3) the current state of the system, and (4) the environmental model, which captures how the state of the environment changes based on system actions. Using a model-predictive control (MPC) approach [6], the evaluator finds an optimal requirement either via strengthening or weakening of the current requirement and produces the corresponding control actions through the planner.
@@ IV-A Environment Model
  • $\mathcal{A}$ is the set of all system actions, $\mathcal{A}=\mathcal{A}_{sys}\cup\mathcal{A}_{env}$, where $\mathcal{A}_{sys}$ is the set of actuator actions and $\mathcal{A}_{env}$ is the set of environmental actions. Note that $\mathcal{A}_{sys}$ and $\mathcal{A}_{env}$ are disjoint, meaning that $\mathcal{A}_{sys}\cap\mathcal{A}_{env}=\emptyset$.
@@
We do not impose restrictions on one particular notation for specifying an environment model, as long as it can be used to generate signals in the format illustrated above. With a more powerful backend like Gurobi [7], one can encode more complex environmental models like non-linear dynamics. For our implementation, we use the MiniZinc modeling language [8], which provides declarative constraints for specifying relationships between different variables of a system.
IV-B Adaptation Trigger
@@
$\exists as_{env}:Seq(\mathcal{A}_{env})\bullet\forall as_{sys}:Seq(\mathcal{A}_{sys})\bullet$
$\quad\quad\forall\textbf{s}:Signal\bullet\textbf{s}{}=\delta^{*}(q_{0},as_{env}\oplus as_{sys})\implies\textbf{s}{}\not\models\varphi$
where $q_{0}$ is the current state of the system and $\oplus$ is an interleaving of two sequences of actions. In other words, the above statement says that the environment can behave in a certain way (through some sequence of events, $a_{env}$) such that where no matter how the system responds, it is unable to satisfy the desired property $\varphi$. The idea is then to carry out degradation to find and satisfy a weaker version of $\varphi$.
@@
$\forall as_{env}:Seq(\mathcal{A}_{env})\bullet\exists as_{sys}:Seq(\mathcal{A}_{sys})\bullet$
$\quad\quad\forall\textbf{s}:Signal\bullet\textbf{s}=\delta^{*}(q_{0},as_{env}\oplus as_{sys})\land\textbf{s}\models\varphi$
@@
Evaluating the above conditions involves generating a predictive signal (s) that describes how the environment evolves given a sequence of system actions. However, carrying this out frequently may incur significant runtime overhead and possibly interfere with the system operations. Thus, instead of evaluating these conditions, our event detector looks for designated degradation and restoration events and uses these as triggers for degradation and recovery, respectively. Degradation events are abnormal events that occur due to an unexpected change in the environment, such as a thruster failure or unusually low water visibility. On the other hand, a restoration event indicates the environment returning to its previous state, such as an improvement in visibility or recovery of an engine thruster.
IV-C Adaptation Problems
@@
Graceful Degradation Problem: Given degradation event $a_{0}$, current requirement $\varphi_{curr}$, find $\varphi_{curr}^{\prime}$ and $as_{sys}$ such that
$\exists\textbf{s}{}_{pred}\cdot(\textbf{s}_{curr}~{}^{\frown}\textbf{s}{}_{pred})$ $\models\varphi_{curr}^{\prime}~{}\land$ (2)
@@
$\forall\textbf{s}{}\bullet\textbf{s}{}\models\varphi_{curr}$ $\Rightarrow\textbf{s}{}\models\varphi_{curr}^{\prime}~{}\land$ (3)
$\textbf{s}{}_{pred}$ $=\delta^{*}(\textbf{s}{}_{curr},\langle a_{0}\rangle~{}^{\frown}as_{sys})~{}\land$ (4)
$a_{0}$ $\in\mathcal{A}_{degrade}$ (5)
Note that the ${}^{\frown}$ is the concatenation operator that is used to link two sequences together (i.e. $\langle s_{1},s_{2}\rangle^{\frown}\langle s_{3},s_{4}\rangle$ results in $\langle s_{1},s_{2},s_{3},s_{4}\rangle$). Degradation actions $\mathcal{A}_{degrade}$ is a subset of all environmental actions $\mathcal{A}_{env}$ (i.e., $\mathcal{A}_{degrade}\subseteq\mathcal{A}_{env}$) defined in Section IV. $\textbf{s}{}_{pred}$ represents the future signal generated by the sequence of control actions $as_{sys}$, and $\textbf{s}{}_{curr}$ represents the signal that encompasses the current state, which is monitored by the sensor. Note also that the new requirement needs to be weaker than the current requirement to achieve degradation, using the definition in Eq. (3), which will be formally defined in Section V.
Informally, this problem involves given a degradation event, how to find control actions that can satisfy an alternative requirement such that it still provides an acceptable level of system utility. We assume that the degradation event can lead the system to violate the current requirement.
Problem IV.2
Recovery Problem: Given restoration event a0superscriptsubscriptnormal-$a_{0}^{\prime}$, current requirement $\varphi_{curr}$, find $\varphi_{curr}^{\prime\prime}$ and $as_{sys}^{\prime}$ such that
@@
$\exists\textbf{s}{}_{pred}\cdot(\textbf{s}_{curr}~{}^{\frown}\textbf{s}{}_{pred})$ $\models\varphi_{curr}^{\prime\prime}~{}\land$ (6)
$\forall\textbf{s}{}\bullet\textbf{s}{}\models\varphi_{curr}^{\prime\prime}$ $\Rightarrow\textbf{s}{}\models\varphi_{curr}~{}\land$ (7)
$\textbf{s}{}_{pred}$ $=\delta^{*}(\textbf{s}{}_{curr},\langle a_{0}^{\prime}\rangle~{}^{\frown}as_{sys}^{\prime})~{}\land$ (8)
$a_{0}$ $\in\mathcal{A}_{restoration}$ (9)
@@
Note that similar to the previous problem definition, the set of restoration actions $\mathcal{A}_{restoration}$ is a subset of all environmental actions $\mathcal{A}_{env}$ ($\mathcal{A}_{restoration}\subseteq\mathcal{A}_{env}$). $\textbf{s}{}_{pred}$ represents the future signal generated by the sequence of control actions $as_{sys}^{\prime}$. Note that in the case of recovery, the new requirement needs to be stronger than the current requirement to achieve recovery, as defined in Eq. (7). The formalism for strengthening will be formally defined in Section V.
@@
The recovery problem is framed as given a restoration event, how to find control actions that can satisfy an alternative requirement such that it can provide a better level of system utility. We assume that the restoration event has the potential to achieve a more desired goal than the current requirement.
V Requirement Weakening and Strengthening
@@
In this section, we present an extension to PSTL to (1) incorporate the concept of requirement weakening and strengthening, and (2) restrict the search space of alternative requirements by providing upper- and lower bounds for the validity domain of PSTL formulas. We also propose metrics to compare multiple PSTL instantiations; these metrics are then used by the MILP solver to find an optimal requirement; i.e., a minimally weakened or maximally strengthened version of the current requirement.
V-A Minimal, Optimal and Current requirements
@@
To enable requirement adaptation, we introduce three new concepts that guide and restrict the range of the requirement space. The minimal requirement, $\varphi_{min}$, represents the lower bound of the PSTL requirement, allowing for the loosening of constraints when necessary for system adaptability. Conversely, the optimal requirement, $\varphi_{opt}$, signifies the upper bound of the PSTL requirement, enabling the strengthening of the reference requirement. We assume that any requirements that are stronger than the optimal requirement do not provide additional utility for the system, and on the contrary, any requirement weaker than the minimal requirement will result in behavior that is deemed unacceptable to the stakeholders or designer.
@@ V-B Extension to PSTL
Suppose $\varphi_{1}$ and $\varphi_{2}$ are two distinct instantiations of the PSTL formula $\varphi$, such that $\varphi_{1}=\varphi(\nu_{1}(\textbf{p}))$, $\varphi_{2}=\varphi(\nu_{2}(\textbf{p}))$, and $\varphi_{1}\neq\varphi_{2}$. We say that $\varphi_{1}$ is weaker than $\varphi_{2}$ if Eq. 11 is true, denoted as $\varphi_{1}\ll\varphi_{2}$; conversely, $\varphi_{1}$ is stronger than $\varphi_{2}$ if Eq. 10 holds, denoted as $\varphi_{1}\gg\varphi_{2}$.
$\forall\textbf{s}\in T\rightarrow D,\textbf{s}{}\models\varphi_{1}\Rightarrow\textbf{s}{}\models\varphi_{2}$ (10)
$\forall\textbf{s}\in T\rightarrow D,\textbf{s}{}\models\varphi_{2}\Rightarrow\textbf{s}{}\models\varphi_{1}$ (11)
@@
Monotonic PSTL
The set of valuations is a partially ordered set with the order operator $\prec$, where $\nu\prec\nu^{\prime}$ iff $\forall i\bullet\nu(\textbf{p})_{i}\leq\nu^{\prime}(\textbf{p})_{i}$ where $\nu(\textbf{p})_{i}$ and $\nu^{\prime}(\textbf{p})_{i}$ are the i-th index of the valuation tuple provided by the valuation function $\nu$ and $\nu^{\prime}$, respectively.
A PSTL formula is called monotonically increasing if Eq. 12 is true, and monotonically decreasing if Eq. 13 is true. A formula is called a monotonic formula if it is either monotonically increasing or decreasing.
$\forall\textbf{s}{}\in T\rightarrow D,\nu\prec\nu^{\prime}\land\textbf{s}{}\models\varphi(\nu(\textbf{p}))\Rightarrow\textbf{s}{}\models\varphi(\nu^{\prime}(\textbf{p}))$ (12)
$\forall\textbf{s}{}\in T\rightarrow D,\nu\prec\nu^{\prime}\land\textbf{s}{}\models\varphi(\nu(\textbf{p}))\Rightarrow\textbf{s}{}\models\varphi(\nu^{\prime}(\textbf{p}))$ (13)
Within the scope of this work, we will only be considering monotonic PSTL formulas, to reduce the search space by eliminating cases where distinct valuations of a PSTL may not necessarily weaken or strengthen the requirements (therefore making the planning strategy more effective). For example, one such requirement is $\varphi~{}\mathcal{U}{}_{[0+\tau,3+\tau]}~{}\psi$, where $\tau$ is a time parameter. In this case, a strict increase or decrease of the parameter $\tau$ does not indicate a weakening or strengthening of the PSTL formula.
@@
The domain validity of a PSTL formula is evaluated with respect to a signal trace s. It is the set of all the parameter values that yield satisfaction starting at time zero for the trace s. We extend the validity domain concept so that it is bounded by minimal and optimal requirements we defined above in Section V-A. Formally, the validity domain of PSTL formula $\varphi$ with respect to a signal s is denoted as $D(\textbf{s}{},\varphi)\subseteq\mathcal{P}$, which consists of a tuple of all values of the parameters $\textbf{v}=(v_{1},...,v_{m})$ within a PSTL formula $\varphi$ such that s satisfies $\varphi(\textbf{v})$. To quantify the inequalities between these valuation tuples, we defined the $\preceq$ operator: $\textbf{v}\preceq\textbf{v}^{\prime}$ if and only if $\forall j\bullet\textbf{v}_{j}\leq\textbf{v}^{\prime}_{j}$. In addition, the $\oplus$ operator in this context is used for time interval concatenation (i.e. $t\oplus[t_{1},t_{2}]$ is equivalent to $[t+t_{1},t+t_{2}]$). Note that to compute the validity domain, we do need an intermediate stage of the form $d(\textbf{s}{},\varphi)\subseteq T\times\mathcal{P}$, which consists of all the tuples $(t,\textbf{v})$ such that $(\textbf{s}{},t)\models\varphi(\textbf{v})$.
@@
$(\textbf{v}_{min}\preceq\textbf{v}{}\preceq\textbf{v}_{opt})$
@@
$\land~{}(\textbf{v}_{min}\preceq\textbf{v}{}\preceq\textbf{v}_{opt})$
Note that $a_{\textbf{v}}$ and $I_{\textbf{v}}$ here either represent constants or the concrete value assignment of the parameter $a$ and $I$ using valuation tuple v. $\textbf{v}{}_{min}$ and $\textbf{v}{}_{opt}$ correspond to the valuations that instantiate the minimal and optimal requirements $\varphi_{min}$ and $\varphi_{opt}$, respectively.
@@
To quantitatively measure and compare the relative restrictiveness between two instantiated formulas (i.e. $\varphi(\nu_{1}(\textbf{p}))$ and $\varphi(\nu_{2}(\textbf{p}))$, which will be abbreviated to $\varphi_{1}$ and $\varphi_{2}$ below) from a PSTL formula, $\varphi$, we introduce two metrics: degree of weakening (Eq. 14), and its inverse metric, degree of strengthening (Eq. 15). These metrics, in turn, based on the robustness of satisfaction in STL, are as follows:
$\begin{split}\Delta_{weak}(\varphi_{1},\varphi_{2},\textbf{s}{},t)&=\rho(\varphi_{2},\textbf{s}{},t)-\rho(\varphi_{1},\textbf{s}{},t)\end{split}$ (14)
$\begin{split}\Delta_{strong}(\varphi_{1},\varphi_{2},\textbf{s}{},t)&=\rho(\varphi_{1},\textbf{s}{},t)-\rho(\varphi_{2},\textbf{s}{},t)\end{split}$ (15)
@@
Recall the example in Section II, with the requirement for the thruster feature: The thruster should provide 100N of thrust within the next second. The parameter 100N is subject to change. This requirement can be formalized in PSTL as $\varphi\equiv\Box_{[0,1]}(\texttt{thrust}>p_{1})$. The original requirement is $\varphi_{origin}\equiv\varphi(p_{1}\mapsto 100)\equiv\Box_{[0,1]}(\texttt{thrust}>100)$, while one possible weakened version is $\varphi_{weak}\equiv\varphi(p_{1}\mapsto 70)\equiv\Box_{[0,1]}(\texttt{thrust}>70)$.
Suppose the system evolves to generate signal s with a thrust of $110$ and $80$, at time 00 and $1$ second, respectively; then $\rho(\varphi_{origin},\textbf{s}{},0)=-20$ (meaning $\varphi_{origin}$ is violated), while $\rho(\varphi_{weak},\textbf{s}{},0)=10$ (meaning $\varphi_{weak}$ is satisfied. Thus, the degree of weakening is measured as $\Delta_{weak}(\varphi_{origin},\varphi_{weak},\textbf{s}{},0)=30$ (meaning requirement weakening from $\varphi_{weak}$ to $\varphi_{origin}$), while degree of strengthening is measured as $\Delta_{strong}(\varphi_{weak},\varphi_{origin},\textbf{s}{},0)=30$ (meaning strengthening from $\varphi_{weak}$ to $\varphi_{origin}$.
V-C Runtime Adaptation as a MILP Formulation

### 08 Apr 2024 v2

@@ 4. Runtime Adaptation Architecture
There are three major components in the proposed adaptation architecture. First, the event detector looks for degradation or restoration events in the environment. Second, the requirement evaluator determines the achievable requirement based on the current environmental conditions. Third, the degradation and recovery planner plan future system actions based on the changing requirements. For example, if visibility decreases below a certain threshold and the weakened requirement states that ”regain contact within the next 5 seconds”, the degradation planner may institute a wait action instead of diving directly or diving with a lower speed.
When activated by the event detector, the requirement evaluator takes various inputs: (1) an environmental event, (2) the current system requirement, (3) the current state of the system, and (4) the environmental model, which captures how the state of the environment changes based on system actions. Using a model-predictive control (MPC) approach (Rawlings and Mayne, 2009), the evaluator finds an optimal requirement either via strengthening or weakening of the current requirement and produces the corresponding control actions through the planner.
@@ 4.1. Environment Model
  • $\mathcal{A}$ is the set of all actions, $\mathcal{A}=\mathcal{A}_{sys}\cup\mathcal{A}_{env}$, where $\mathcal{A}_{sys}$ is the set of actuator actions and $\mathcal{A}_{env}$ is the set of environmental actions. Note that $\mathcal{A}_{sys}$ and $\mathcal{A}_{env}$ are disjoint, meaning that $\mathcal{A}_{sys}\cap\mathcal{A}_{env}=\emptyset$.
@@
We do not impose restrictions on one particular notation for specifying an environment model, as long as it can be used to generate signals in the format illustrated above. With a more powerful backend like Gurobi (Rimmi Anand and Kumar, 2017), one can encode more complex environmental models like non-linear dynamics. For our implementation, we use the MiniZinc modeling language (Nethercote et al., 2007), which provides declarative constraints for specifying relationships between different variables of a system.
4.2. Adaptation Trigger
@@
$\exists\textbf{a}_{env}:Seq(\mathcal{A}_{env})\bullet\forall\textbf{a}_{sys}:Seq(\mathcal{A}_{sys})\bullet$
$\quad\quad\forall\textbf{s}:\textbf{S}\bullet\textbf{s}{}=\delta^{*}(q_{0},\textbf{a}_{env}\oplus\textbf{a}_{sys})\implies(\textbf{s}{},0)\not\models\varphi$
where S is the set of all signals, $q_{0}$ is the current state of the system, $\oplus$ is an interleaving of two sequences of actions. In other words, the above statement says that the environment can behave in a certain way (through some sequence of actions, $\textbf{a}_{env}$) such that no matter how the system responds, it is unable to satisfy the property $\varphi$. The idea is then to carry out degradation to find and satisfy a weaker version of $\varphi$.
@@
$\forall\textbf{a}_{env}:Seq(\mathcal{A}_{env})\bullet\exists\textbf{a}_{sys}:Seq(\mathcal{A}_{sys})\bullet$
$\quad\quad\forall\textbf{s}:\textbf{S}\bullet\textbf{s}=\delta^{*}(q_{0},\textbf{a}_{env}\oplus\textbf{a}_{sys})\land(\textbf{s},0)\models\varphi$
@@
Evaluating the above conditions involves generating a predictive signal (s) that describes how the environment evolves given a sequence of system actions. However, carrying this out frequently may incur significant runtime overhead and possibly interfere with the system operations. Thus, instead of evaluating these conditions, our event detector looks for designated degradation and restoration events ($A_{degrade}$ and $A_{restore}$) and uses these as proxy triggers for degradation and recovery, respectively. Degradation events are abnormal events that occur due to an unexpected change in the environment, such as a thruster failure or unusually low water visibility. On the other hand, a restoration event indicates the environment returning to its previous state, such as an improvement in visibility or recovery of an engine thruster.
4.3. Adaptation Problems
@@
Graceful Degradation Problem. Given degradation event $a_{0}\in\mathcal{A}_{degrade}$ and current requirement $\varphi_{curr}$, find $\varphi_{curr}^{\prime}$ and action sequence a such that
@@
(2) $\forall\textbf{s}{}:\textbf{S}\bullet(\textbf{s}{},0)\models\varphi_{curr}\Rightarrow(\textbf{s}{},0)\models\varphi_{curr}^{\prime}~{}\land$
@@
(3) $\exists\textbf{s}{}_{pred}:\textbf{S}\cdot(\textbf{s}_{curr}~{}^{\frown}\textbf{s}{}_{pred},0)\models\varphi_{curr}^{\prime}~{}\land$
(4) $\qquad\textbf{s}{}_{pred}=\delta^{*}(q_{0},\langle a_{0}\rangle~{}^{\frown}\textbf{a})$
@@
where $q_{0}$ is the current state of the system (encoded in $\textbf{s}_{curr}$) and ${}^{\frown}$ is the concatenation operator that is used to link two sequences together (i.e., $\langle s_{1},s_{2}\rangle^{\frown}\langle s_{3},s_{4}\rangle$ results in $\langle s_{1},s_{2},s_{3},s_{4}\rangle$). Degradation actions $\mathcal{A}_{degrade}$ is a subset of all environmental actions $\mathcal{A}_{env}$ (i.e., $\mathcal{A}_{degrade}\subseteq\mathcal{A}_{env}$) defined in Section 4. $\textbf{s}{}_{pred}$ represents the future signal generated by the sequence of actions a, and $\textbf{s}{}_{curr}$ represents the signal that encompasses the current state, which is monitored by the sensor. Note also that the new requirement needs to be weaker than the current requirement to achieve degradation, using the definition in Eq. (2), which will be formally defined in Section 5.
@@
Informally, this problem involves given a degradation event, how to find control actions that can satisfy an alternative requirement such that it still provides an acceptable level of system utility.
Problem 4.2 ().
Recovery Problem. Given restoration event $a_{0}\in\mathcal{A}_{restore}$ and current requirement $\varphi_{curr}$, find $\varphi_{curr}^{\prime\prime}$ and a such that
(5) $\forall\textbf{s}{}:\textbf{S}\bullet(\textbf{s}{},0)\models\varphi_{curr}^{\prime\prime}\Rightarrow(\textbf{s}{},0)\models\varphi_{curr}~{}\land$
(6) $\exists\textbf{s}{}_{pred}:\textbf{S}\cdot(\textbf{s}_{curr}~{}^{\frown}\textbf{s}{}_{pred},0)\models\varphi_{curr}^{\prime\prime}~{}\land$
(7) $\qquad\textbf{s}{}_{pred}=\delta^{*}(q_{0},\langle a_{0}\rangle~{}^{\frown}\textbf{a})$
Note that in the case of recovery, the new requirement needs to be stronger than the current requirement to achieve recovery, as defined in Eq. (5). The formalism for strengthening will be formally defined in Section 5. Informally, the recovery problem is framed as given a restoration event, how to find control actions that can satisfy an alternative requirement such that it can provide an improved level of system utility.
5. Requirement Weakening and Strengthening
@@
In this section, we present an extension to PSTL to (1) incorporate the concept of requirement weakening and strengthening, and (2) restrict the search space of alternative requirements by providing upper and lower bounds for the validity domain of PSTL formulas. We also propose metrics to compare multiple PSTL instantiations. These metrics are then used by the MILP solver to find an optimal requirement; i.e., a minimally weakened or maximally strengthened version of the current requirement.
5.1. Minimal, Optimal and Current requirements
@@
To enable requirement adaptation, we introduce three new concepts that guide and restrict the range of the requirement space. The minimal requirement, $\varphi_{min}$, represents the lower bound of the PSTL requirement, allowing for the loosening of constraints when necessary for system adaptability. Conversely, the optimal requirement, $\varphi_{opt}$, signifies the upper bound of the PSTL requirement, enabling the strengthening of the reference requirement. We assume that any requirements that are stronger than the optimal requirement do not provide additional utility for the system, and on the contrary, any requirement weaker than the minimal requirement will result in behavior that is deemed unacceptable to stakeholders.
@@ 5.2. Extension to PSTL
Suppose $\varphi_{1}$ and $\varphi_{2}$ are two distinct instantiations of the PSTL formula $\varphi$, such that $\varphi_{1}=\varphi(\nu_{1}(\textbf{p}))$, $\varphi_{2}=\varphi(\nu_{2}(\textbf{p}))$, and $\varphi_{1}\neq\varphi_{2}$. We say that $\varphi_{1}$ is weaker than $\varphi_{2}$ if and only if Eq. 9 is true, denoted as $\varphi_{1}\ll\varphi_{2}$; conversely, $\varphi_{1}$ is stronger than $\varphi_{2}$ if and only if Eq. 8 holds, denoted as $\varphi_{1}\gg\varphi_{2}$.
(8) $\forall\textbf{s}\in T\rightarrow D\cdot(\textbf{s}{},0)\models\varphi_{1}\Rightarrow(\textbf{s}{},0)\models\varphi_{2}$
(9) $\forall\textbf{s}\in T\rightarrow D\cdot(\textbf{s}{},0)\models\varphi_{2}\Rightarrow(\textbf{s}{},0)\models\varphi_{1}$
@@
The validity domain of a PSTL formula (Asarin et al., 2012), evaluated with respect to signal trace s, is the set of all the parameter values that yield satisfaction for the trace s. We extend the validity domain concept so that it is bounded by minimal and optimal requirements that we defined above in Section 5.1.
We first define the $\preceq$ operator: $\textbf{v}\preceq\textbf{v}^{\prime}$ if and only if $\forall j\bullet\textbf{v}_{j}\leq\textbf{v}^{\prime}_{j}$. The $\oplus$ operator in this context is used for time interval concatenation (i.e., $t\oplus[t_{1},t_{2}]$ is equivalent to $[t+t_{1},t+t_{2}]$). Additionally, $\textbf{v}{}_{min}$ and $\textbf{v}{}_{opt}$ correspond to the valuations that instantiate the minimal and optimal requirements $\varphi_{min}$ and $\varphi_{opt}$, respectively. Then, the validity domain of PSTL formula $\varphi$ with respect to a signal s, bounded by $\textbf{v}{}_{min}$ and $\textbf{v}{}_{opt}$, is denoted as $d(\textbf{s}{},\varphi)$ and defined in the following way:
@@
$(\textbf{v}_{min}\preceq\textbf{v}{}\preceq\textbf{v}_{opt})\}$
@@
$\land~{}(\textbf{v}_{min}\preceq\textbf{v}{}\preceq\textbf{v}_{opt})\}$
Note that $a_{\textbf{v}}$ and $I_{\textbf{v}}$ here either represent constants or the concrete value assignment of the parameter $a$ and $I$ using valuation tuple v. The resulting set, $d(\textbf{s},\varphi)$, is a set of all 2-element tuples of form $(t,\nu(\textbf{p}))$, where $(\textbf{s},t)\models\varphi(\nu(\textbf{p}))$.
@@
To quantitatively measure and compare the relative restrictiveness between two instantiated formulas (i.e., $\varphi(\nu_{1}(\textbf{p}))$ and $\varphi(\nu_{2}(\textbf{p}))$, which will be abbreviated to $\varphi_{1}$ and $\varphi_{2}$ below), we introduce two metrics: degree of weakening (Eq. 10), and its inverse metric, degree of strengthening (Eq. 11). These metrics are defined based on the robustness of satisfaction, as follows:
(10) $\begin{split}\Delta_{weak}(\varphi_{1},\varphi_{2},\textbf{s}{},t)&=\rho(\varphi_{2},\textbf{s}{},t)-\rho(\varphi_{1},\textbf{s}{},t)\end{split}$
(11) $\begin{split}\Delta_{strong}(\varphi_{1},\varphi_{2},\textbf{s}{},t)&=\rho(\varphi_{1},\textbf{s}{},t)-\rho(\varphi_{2},\textbf{s}{},t)\end{split}$
@@
Recall the example in Section 2, with the requirement for the thruster feature: The thruster should provide 100N of thrust within the next second. The parameter 100N is subject to change. This requirement can be formalized in PSTL as $\varphi\equiv\Box_{[0,1]}(\texttt{thrust}>p_{1})$. The original requirement is $\varphi_{origin}\equiv\varphi(p_{1}\mapsto 100)\equiv\Box_{[0,1]}(\texttt{thrust}>100)$, while one possible weakened version is $\varphi_{weak}\equiv\varphi(p_{1}\mapsto 70)\equiv\Box_{[0,1]}(\texttt{thrust}>70)$.
Suppose the system evolves to generate signal s with a thrust of $110$ and $80$, at time 00 and $1$ second, respectively; then $\rho(\varphi_{origin},\textbf{s}{},0)=-20$ (meaning $\varphi_{origin}$ is violated), while $\rho(\varphi_{weak},\textbf{s}{},0)=10$ (meaning $\varphi_{weak}$ is satisfied. Thus, the degree of weakening is measured as $\Delta_{weak}(\varphi_{origin},\varphi_{weak},\textbf{s}{},0)=30$ (meaning requirement weakening from $\varphi_{weak}$ to $\varphi_{origin}$), while degree of strengthening is measured as $\Delta_{strong}(\varphi_{weak},\varphi_{origin},\textbf{s}{},0)=30$ (meaning strengthening from $\varphi_{weak}$ to $\varphi_{origin}$).
5.3. Runtime Adaptation as a MILP Formulation